OpenStudy (anonymous):

Calculus: Find the critical numbers for f(x) = 2sec(x)+tan(x) on the interval 0<=x<2pi (0<=360) I'm stuck x.x

OpenStudy (anonymous):

how far have you gotten so far?

OpenStudy (anonymous):

All I really did was rewrite the equation as f(x)= 2/cosx + sinx/cosx What should I do next? sec having critical numbers at multiples of pi/2, tan also??

OpenStudy (loser66):

what do you get for f'(x)? how to get critical points from that f'(x) show your work, please

OpenStudy (anonymous):

actually.. using the original form was easier, f'(x) = 2secxtanx+secx^2 but then how would I get the critical numbers? I

OpenStudy (loser66):

factor sec out,

OpenStudy (loser66):

you get sec (2tanx +sec) =0 so, sec =0 or 2tan +sec =0 however, sec =1/cos therefore, it never be 0, so you just have 2tan +sec =0. it is \[2\frac{sin}{cos}+\frac {1}{cos}=\frac {2sin+1}{cos}\]it is =0 iff 2sinx +1 =0 , sinx = -1/2 x = ??? that is your critical points. plug them ( if there are more than 1 solution) into original function to get value of critical points. that's all you need

OpenStudy (loser66):

tag other if you have more question, please. Sorry for that. I want to eat something. hehe

OpenStudy (anonymous):

Ah, thanks! I got 3/sqrt(3) and -3/sqrt(3) as my final answer. is it correct?